Vitamin C or Ascorbic acid (C6H8O6) is a water soluble vitamin. A solution is prepared by dissolving 80.5 g of Ascorbic acid in 210 g of water. The resulting solution has a density of 1.22 g/ml. Calculate the concentration of Ascorbic acid in terms of

a) Mass Percent    b) Mole fraction    c) Molality        d) Molarity
Answer #1

Mass percent =80.5 / 210 * 100

Moles ascorbic acid == 80.5g / 176.2g/mol =0.45moles

Moles water = 210/18 =11.67 moles

Mole fraction =0.45 / (11.67 + 0.45)

Molality = 0.45 moles ascorbic acid / (80.5 + 210 ) g solution=0.00154m

Volume of solution = mass / density = 290.5 g / 1.22g/ml =238ml

Molarity = 0.45 moles / 0.238 L =1.89M